OpenStudy (anonymous):

How do I solve this math problem? Please don't give me the answer. 3+p/3 - 4p=1 - p+7/2

OpenStudy (loser66):

still need help?

OpenStudy (anonymous):

Yes please!!!!!!

OpenStudy (loser66):

\(\dfrac{3+p}{3}-4p =1-\dfrac{p+7}{2}\), right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

When I multiply and then divide a term by the same number, that term doesn't change, right? so, if I do something like this, can you get it? (Nothing change) \(\dfrac{3+p}{3}\color{blue}{\dfrac{2}{2}}-4p\color{red}{\dfrac{6}{6}} =1\color{green}{\dfrac{6}{6}}-\dfrac{p+7}{2}\color{purple}{\dfrac{3}{3}}\)

OpenStudy (anonymous):

I kind of get what you are saying

OpenStudy (anonymous):

But where did you get those numbers from?

OpenStudy (loser66):

I did it to get the same denominator for all terms. It comes from the available denominator 3 and 2, then the same denominator is 6. Therefore, the first term \(\dfrac{3+p}{3}\) need 2 to get 6 (when multiplying)

OpenStudy (anonymous):

That makes more sense

OpenStudy (loser66):

the second term is 4p, hence the denominator is 1, then we need *6 to get common denominator, ok?

OpenStudy (anonymous):

Ok

OpenStudy (loser66):

the same reason with the right hand side, ok?

OpenStudy (anonymous):

Ok

OpenStudy (loser66):

Now, all are the same denominator, we just distribute and combine like terms: \(\dfrac{(3+p)2}{6}-\dfrac{6(4p)}{6} =\dfrac{6}{6}-\dfrac{(p+7)3}{6}\) ok?

OpenStudy (loser66):

\(\dfrac{6+2p-24p}{6}=\dfrac{6-(3p+21)}{6}\\\dfrac{6-22p}{6}=\dfrac{3p-15}{6}\) got me so far?

OpenStudy (loser66):

Not finish yet, but you need understand before moving on. Just few more steps and we are done.

OpenStudy (anonymous):

Yes I understand

OpenStudy (anonymous):

Sorry for the late reply